On their 1200 km-trip to Baguio, the Esperas first took a plane and later a van. The van, traveling at 48 kph took 2 hrs. longer than the plane, traveling at 240 kph. How long did the whole trip take?


Pls. ans. this...

Answer by Jk22(389)

Let D=1200km the distance to be covered,

tv the time taken by van travel
tp the one needed for plane travel.

Then the searched quantity is tv+tp, knowing that : tv=2+tp in hours

The distance is 48*tv+240*tp by taking the respective speeds, and this should be equal to D, hence,

48*(tp+2) + 240*tp = 1200

288*tp + 96 = 1200

288*tp = 1104

tp = 552/144 = 23/6h, the total time is : tp+tv=2*tp+2=23/3+2=29/3=9+2/3 hours

or 9hours and 40minutes total travel time.